# Benchmarking the performance of quantum computing software

### v1

@@ Introduction
The promise of quantum computation lies in its ability to perform specific tasks more efficiently than otherwise possible using classical methods alone. However, quantum computers do not operate in isolation and require classical computing resources for data pre- and post-processing. As quantum computers continue to grow in size, it is imperative that the associated classical computing costs be evaluated for scalability, and the construction, manipulation, and optimization of quantum circuits play an out-sized role in terms of classical resource overhead. Moreover, the performance of quantum computation software for output circuit quality, run-time, and memory consumption is critical to successful adoption of this technology. There is a variety of quantum software development kits (SDKs) that perform all or some fraction of quantum circuit pre-execution processing, such as Braket [1], BQSKit [2], Cirq [3], Qiskit [4], Qiskit Transpiler Service (QTS) extension [5], Staq [6], and Tket [7], amongst others. Here we define an SDK to be a collection of software development tools within a single installable package. The need to evaluate the performance of these software stacks has led to the creation of several collections of assessment circuits [8, 9] and Hamiltonians [10], as well as some examples of how to execute collections of circuits in a benchmarking framework [11, 12].
@@
To date, there has been no systematic study of the relative performance of mainstream quantum computing SDKs over extensive collections of quantum circuits at scale, nor has there been a flexible method by which these comparisons can be easily generated. Such a study is critical as quantum devices and experiments push towards 100-qubits or more [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32], and the differences in software performance and scaling become pronounced. An open-source test suite that not only evaluates multiple SDKs at these scales but also provides an execution framework that yields uniform testing over quantum software with disparate feature sets and capabilities would help researchers, developers, and end users alike ascertain the relative value of quantum computing software packages when targeting current and future quantum computing devices.
@@ I.1 Circuit construction and manipulation
Although nominally a tiny part of an overall workflow time budget, as compared to circuit transpilation, measuring the timing of circuit construction and manipulation gives a holistic view of quantum SDK performance. Moreover, if suitably chosen, such tests can provide insights into other parts of the entire circuit compilation process. The present version of Benchpress includes 12 such tests, with tests aimed at representing scenarios encountered during real-world SDK usage. Circuit construction includes eight tests that look at timing information needed to build 100-qubit circuits for families of circuits such as Quantum Volume (QV) [36] (test_QV100_build), Hamiltonian simulation [15] (test_DTC100_set_build), random Clifford circuits [37] (test_clifford_build), 16-qubit iterative construction of multi-controlled gates (test_multi_control_circuit), and parameterized ansatz circuits with circular entangling topology (test_param_circSU2_100_build). We will reuse many of these circuits in device benchmarking. We also include the time to bind values to parameterized circuits (test_param_circSU2_100_bind) and import from OpenQASM files into the construction category. This latter set of QASM tests includes importing a 100-qubit QV circuit and reading a file with an integer corresponding to a 301-bit classical register, test_QV100_qasm2_import and test_bigint_qasm2_import respectively. Our focus on $100$-qubit circuits stems from the need for sufficient complexity for gathering faithful timing information, and the fact that these circuits are within the number of qubits available on present-day quantum processors.
@@
For circuit manipulation, both Qiskit and Tket complete all tests, with Qiskit completing all four tests in $5.5$ seconds versus $7.1$ for Tket. The large number of skipped tests for Braket is due to the lack of basis transformation capabilities. However, add-on packages allow this to be done in other SDKs and then passed back to Braket [41]. The two failed tests for Cirq are due to using the basis set of gates: $RZ$, $X$, $\sqrt{X}$, and $CZ$, which despite being supported gates, hits a recursion limit. In addition to timing information, the multi-controlled decomposition test allows for examining the quality of the synthesis algorithms in each SDK via the number of two-qubit gates in the output circuit. Here, Tket produced the circuit with the fewest two-qubit gates, $4457$, versus $7349$ and $17414$ for Qiskit and Cirq, respectively.
@@ I.2 Device and abstract transpilation
Due to their vast array of possible input parameters and a large fraction of overall runtime, transpilation tests form the bulk of the tests in Benchpress. We split these tests into two groups depending on whether they target a model of a real quantum device, or if their target is an abstract topology defined by a generating function. We label these as “device" and “abstract" transpilation tests, respectively. These tests differ because device testing targets a fixed model of a quantum system, regardless of input quantum circuit size, and includes error rates that can be utilized in noise-aware compilation routines. In contrast, abstract transpilation tests take an input circuit and finds the smallest topology compatible with the circuit. In this manner, we can benchmark SDKs across arbitrary circuit sizes and topology families, and allowing for user configuration of the basis gates in the default.conf file, see Sec. (III).
@@
Figure 3 shows the results for the five SDKs that support transpilation, see Fig. (1), over the combined set of device and abstract tests. Out of $1054$ total tests, $22$ are device transpilation tests larger than the target Heron device and are SKIPPED regardless of the SDK. In addition, the Staq compiler takes OpenQASM files as input and thus is unable to execute the Hamiltonian simulation tests, defaulting to those tests being skipped. Only passing tests are shown in Fig. (3). As Qiskit is the only SDK that passes the entire set of transpilation tests, we use those results as a baseline when presenting results. In what follows, we focus on the statistics of ensembles of circuits as a whole, as opposed to looking at individual tests cases. A lower-level analysis can be done using the published results [Sec. (IV)].
@@ II Discussion
In the end, this work aims to benefit the community of quantum researchers, developers, and end-users at large by providing a trusted source by which absolute and/or relative improvements in quantum software can be faithfully examined. For researchers and developers this can mean highlighting performance bottlenecks that need to be improved. For example, the small number of circuits on which both Tket and QTS outperform Qiskit in terms of 2Q-gate depth are of interest to us, and identify areas for improved circuit synthesis and routing, respectively in Qiskit. For others, this work may serve as a guide for where to target future SDK improvements for improving runtime and/or fixing coding bugs. Finally, end-users can use this work as a guide for selecting the appropriate SDK, or perhaps combinations of SDKs, that are optimal for a given task. In all cases, shedding light on the performance of quantum computing software can only push the field further and aid in the successful adoption of this computing paradigm.
@@ Acknowledgements.
The authors thank Julien Gacon, Alexander Ivrii, and Jake Lishman for helpful discussions.

### v2

@@ Introduction
The promise of quantum computation lies in its ability to perform specific tasks more efficiently than otherwise possible using classical methods alone. However, quantum computers do not operate in isolation and require classical computing resources for data pre- and post-processing. As quantum computers continue to grow in size, it is imperative that the associated classical computing costs be evaluated for scalability, and the construction, manipulation, and optimization of quantum circuits play an out-sized role in terms of classical resource overhead. Moreover, the performance of quantum computation software for output circuit quality, run-time, and memory consumption is critical to successful adoption of this technology. There is a variety of quantum software development kits (SDKs) that perform all or some fraction of quantum circuit pre-execution processing, such as Braket [1], BQSKit [2], Cirq [3], Qiskit [4], Qiskit Transpiler Service (QTS) extension [5], Staq [6], and Tket [7], amongst others, that focus primarily on quantum circuit construction, manipulation, and optimization. Here we define an SDK to be a collection of software development tools within a single installable package. The need to evaluate the performance of these software stacks has led to the creation of several collections of assessment circuits [8, 9, 10] and Hamiltonians [11], as well as some examples of how to execute collections of circuits in a benchmarking framework [12, 13].
@@
To date, there has been no systematic study of the relative performance of mainstream quantum computing SDKs over extensive collections of quantum circuits at scale, nor has there been a flexible method by which these comparisons can be easily generated. Such a study is critical as quantum devices and experiments push towards 100-qubits or more [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35], and the differences in software performance and scaling become pronounced. An open-source test suite that not only evaluates multiple SDKs at these scales but also provides an execution framework that yields uniform testing over quantum software with disparate feature sets and capabilities would help researchers, developers, and end users alike ascertain the relative value of quantum computing software packages when targeting current and future quantum computing devices. With fault-tolerant computation out of reach in the near-term, this initial version of Benchpress is focused on test cases and metrics compatible with execution on noisy quantum processors.
@@ I.1 Circuit construction and manipulation
Although nominally a tiny part of an overall circuit processing workflow time budget, as compared to circuit transpilation, measuring the timing of circuit construction and manipulation gives a holistic view of quantum SDK performance. Moreover, if suitably chosen, such tests can provide insights into other parts of the entire circuit compilation process. The present version of Benchpress includes 12 such tests, with tests aimed at representing scenarios encountered during real-world SDK usage. Circuit construction includes eight tests that look at timing information needed to build 100-qubit circuits for families of circuits such as Quantum Volume (QV) [39] (test_QV100_build), Hamiltonian simulation [16] (test_DTC100_set_build), random Clifford circuits [40] (test_clifford_build), 16-qubit iterative construction of multi-controlled gates (test_multi_control_circuit), and parameterized ansatz circuits with circular entangling topology (test_param_circSU2_100_build). We will reuse many of these circuits in device benchmarking. We also include the time to bind values to parameterized circuits (test_param_circSU2_100_bind) and import from OpenQASM files into the construction category. This latter set of QASM tests includes importing a 100-qubit QV circuit and reading a file with an integer corresponding to a 301-bit classical register, test_QV100_qasm2_import and test_bigint_qasm2_import respectively. Our focus on $100$-qubit circuits stems from the need for sufficient complexity for gathering faithful timing information, and the fact that these circuits are within the number of qubits available on present-day quantum processors.
@@
For circuit manipulation, both Qiskit and Tket complete all tests, with Qiskit completing all four tests in $5.5$ seconds versus $7.1$ for Tket. The large number of skipped tests for Braket is due to the lack of basis transformation capabilities. However, add-on packages allow this to be done in other SDKs and then passed back to Braket [44]. The two failed tests for Cirq are due to using the basis set of gates: $RZ$, $X$, $\sqrt{X}$, and $CZ$, which despite being supported gates, hits a recursion limit. In addition to timing information, the multi-controlled decomposition test allows for examining the quality of the synthesis algorithms in each SDK via the number of two-qubit gates in the output circuit. Here, Tket produced the circuit with the fewest two-qubit gates, $4457$, versus $7349$ and $17414$ for Qiskit and Cirq, respectively. The relative gate counts and depth for the circuit manipulation tests are shown in Supplementary Fig. (LABEL:supp-fig:stats).
@@ I.2 Device and abstract transpilation
Due to their vast array of possible input parameters and a large fraction of overall runtime, transpilation tests form the bulk of the tests in Benchpress. We split these tests into two groups depending on whether they target a model of a real quantum device, or if their target is an abstract topology defined by a generating function. We label these as “device" and “abstract" transpilation tests, respectively. These tests differ because device testing targets a fixed model of a quantum system, regardless of input quantum circuit size, and includes error rates that can be utilized in noise-aware compilation routines. Noise aware heuristics can have an impact on both the duration of the compilation process, as well as the two-qubit gate count and depth; they can paradoxically lead to worsened performance if applied overly aggressively. However, because we do not execute the resulting circuits on hardware, the impact of these techniques on output fidelity is not included. In contrast, abstract transpilation tests take an input circuit and finds the smallest topology compatible with the circuit. In this manner, we can benchmark SDKs across arbitrary circuit sizes and topology families, and allowing for user configuration of the basis gates in the default.conf file, see Sec. (III).
@@
Figure 3 shows the results for the five SDKs that support transpilation, see Fig. (1), over the combined set of device and abstract tests. Out of $1054$ total tests, $22$ are device transpilation tests larger than the target Heron device and are SKIPPED regardless of the SDK. In addition, the Staq compiler takes OpenQASM files as input and thus is unable to execute the Hamiltonian simulation tests, defaulting to those tests being skipped. Only passing tests are shown in Fig. (3). As Qiskit is the only SDK that passes the entire set of transpilation tests, we use those results as a baseline when presenting results. In what follows, we focus on the statistics of ensembles of circuits as a whole, as opposed to looking at individual tests cases. A lower-level analysis can be done using the published results [Sec. (IV)]. A break down of the test results for each of the open-source test libraries used here is presented in the Supplemental Material Sec. (LABEL:supp-sec:groupings).
@@
Because we target a fake IBM Quantum system that includes timing information, we can use the device transpilation tests to gauge the relative timing for circuit transpilation verses that of actual execution on hardware. To this end, we schedule each Qiskit circuit to obtain the execution time on chip, and add the idle time between circuits, called the default_rep_delay in Qiskit, to get the total time per circuit execution  [^5] This data was added after the initial version of Benchpress, and is not included in the published results.. Multiplying this value by the default number of executions for IBM Quantum hardware, currently $4096$, gives us a good estimate for circuit execution time without requiring hardware usage. For the Qiskit device transpilation tests, the total compilation time was $1055$ sec, whereas the total execution time of all circuits would be $968$ sec. Restricting ourselves to circuits with $\geq 100$ qubits the total compilation time is $\sim 1.4x$ longer than the $78$ sec hardware runtime; the compilation costs begin to dominate at larger qubit counts. Table (2) shows that other comparable SDKs such as BQSKit and Tket take one- to two-orders of magnitude longer to compile circuits, while only marginally increasing the overall circuit depth, indicating that the ratio of compilation to run-time is more pronounced when using these compilation stacks to target the same quantum device. Note however that is conclusion is dependent on the targeted quantum hardware modality. For example, platforms such as trapped-ion systems have greater connectivity and, in general, require less compilation time because of this. In addition, trapped-ion run-times are two-orders of magnitude or more longer than those on superconducting devices, e.g. see Ref. ([48]), and therefore the ratio of compilation to hardware run-times may be less than those presented here when targeting diffing hardware platforms.
@@ II Discussion
In the end, this work aims to benefit the community of quantum researchers, developers, and end-users at large by providing a trusted source by which absolute and/or relative improvements in quantum software can be faithfully examined. For researchers and developers this can mean highlighting performance bottlenecks that need to be improved. For example, the small number of circuits on which both Tket and QTS outperform Qiskit in terms of 2Q-gate depth are of interest to us, and identify areas for improved circuit synthesis and routing, respectively in Qiskit. For others, this work may serve as a guide for where to target future SDK improvements for improving runtime and/or fixing coding bugs. As an open-source project, Benchpress is aimed at benefiting the quantum community as a whole, and we welcome contributions in the form of bug fixes, code improvements, and new tests cases at the projects Github website [50]. Finally, end-users can use this work as a guide for selecting the appropriate SDK, or perhaps combinations of SDKs, that are optimal for a given task. In all cases, shedding light on the performance of quantum computing software can only push the field further and aid in the successful adoption of this computing paradigm.
@@ Acknowledgements.
The authors thank Julien Gacon, Alexander Ivrii, and Jake Lishman for helpful discussions. L.B., M.T., and A.J-A were supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Co-design Center for Quantum Advantage (C2QA) under contract number DE-SC0012704.